If 2 cosθ = a +
1
a
, prove that 2 cos 2θ = a
2 +
1
a
2
and
2 cos 3θ = a
3 +
1
a
3
 

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
2 cos theta = a+1/a
squaring both sides we get,(2 cos theta)^2=( a+1/a)^2
=>4 cos^2 theta=a^2+1/(a^2)+2
=>4 cos^2 theta=a^2+1/(a^2)+2
=>4 cos^2 theta-2=a^2+1/(a^2)
=>2(2cos^2 theta-1)=a^2+1/(a^2)
=>2 cos2 theta=a^2+1/(a^2) - hence first part is proved.
 
cubing both sides we get, (2 cos theta)^3=(a+1/a)^3
=>8cos^3 theta=a^3+1/(a^3)+3(a+1/a)
=>8cos^3 theta=a^3+1/(a^3)+3(2 cos theta) [from ques.]
=>8cos^3 theta-6 cos theta= a^3+1/(a^3)
=>2(4cos^3 theta - 3cos theta)=a^1/3 +1/(a^3)
=>2 cos3 theta =a^1/3 +1/(a^3) - hence second part is proved.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)